OpenStudy (anonymous):

a garden has a perimeter of 92 feet. If the length of the garden is one foot greater than twice its width, what is the length of the garden, in feet?

OpenStudy (anonymous):

let the width be 'x' and length be 2x+1 \[perimetre=2(l+w)\] \[92=2(2x+1+x)\] \[92/2=3x+1\] \[46=3x+1\] \[3x=46-1\] \[x=45/3=15\] so width=15 feet and length=2(15)+1=30+1=31 feet

OpenStudy (anonymous):

thanks a lot

OpenStudy (anonymous):

always welcome :)